OpenStudy (anonymous):

Henry drove along a straight highway at a constant speed of 60 miles per hour. He observed a pink car one-half mile behind him. The pink car was moving at an average speed that allowed it to pull up even with Henry in exactly three minutes. What was the pink car's average speed, in miles per hour?

OpenStudy (anonymous):

Well what speed difference would allow the car to close a gap of half a mile in 3 minutes?

OpenStudy (anonymous):

just using basic r * t = d

OpenStudy (anonymous):

hello?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so the answere is 70mph?

OpenStudy (anonymous):

yep!

OpenStudy (anonymous):

ktklown is correct with 70mph, the solution for r in the equation below:\[r* \frac{3 }{60}\text{=} \frac{1}{2}+60\text{ }\frac{3 }{60}\]